### v5

# Inverse Reinforcement Learning with Dynamic Reward Scaling for LLM Alignment

Abstract
Alignment is vital for safely deploying large language models (LLMs). Existing techniques are either reward-based–train a reward model on preference pairs and optimize with reinforcement learning (RL)–or reward-free–directly fine-tune on ranked outputs. Recent research show that well-tuned reward-based pipelines remain the most robust, and single-response demonstrations can outperform pairwise preference data. However, there still exist two key challenges: (1) imbalanced safety dataset that overrepresent common hazards while neglecting long-tail threats; and (2) static reward models that ignore task difficulty, limiting optimization efficiency and attainable gains. To address these limitations, we propose DR-IRL, which Dynamically adjusts Rewards through Inverse Reinforcement Learning. We first train category‑specific reward models using a balanced safety dataset of seven harmful categories as demonstration via IRL. Then we enhance Group Relative Policy Optimization (GRPO) by introducing dynamic reward scaling–adjusting rewards by task difficulty–data-level hardness by text encoder cosine similarity, model-level responsiveness by reward gaps. Extensive experiments across various benchmarks and LLMs demonstrate that DR-IRL outperforms all baseline methods in safety alignment while maintaining usefulness.
1 Introduction
Large language models (LLMs) are prone to generating harmful responses when faced with malicious queries [5, 16], especially jailbreak attacks [61] like adversarial suffixes [63, 6] or carefully crafted disguises [45]. Defensive techniques like representation engineering [64], machine unlearning [25], and safeguarding [47] have been proposed, but their reliance on external intervention limits their applicability. Consequently, aligning LLMs is crucial for ensuring their reliability in real-world applications [28].
Current LLM alignment approaches [10] fall into two streams. Reward-based pipelines first fit a reward model to preference data (input paired with two responses, one preferred) [32] and then optimize the model with reinforcement learning (RL) (e.g., PPO [37]) using that signal [7, 39, 2]. While reward-free pipelines skip the reward model and fine-tune LLMs directly on ranked responses [55, 15] (e.g., DPO [34]). Recently, Xu et al. [52] observed that reward-free methods tend to fail when preference data deviate from base model outputs and perform poorly on challenging tasks such as code generation, while well-tuned reward-based pipelines remain robust across diverse benchmarks. Building on this, Li et al. [22] demonstrate that reward learning from demonstration data (input paired with a single example response) can outperform approaches based on preference data, which are costly to collect and still fail to fully capture human values [3, 20].
Despite the advantages, there still exist two key limitations. (1) The training dataset is typically oversampled on certain categories and thus ignore long-tail, high-risk threats [51], underscoring the need for balanced safety datasets. (2) Traditional reward models are static: their fixed reward ignore task difficulty, limiting optimization efficiency and potential performance gains [29, 42].
To address these limitations, we propose DR-IRL, an approach which Dynamically adjusts Reward based on task difficulty via Inverse Reinforcement Learning. Inspired by Li et al. [22], we replace costly preference labels with demonstration data, a balanced Chain of Draft (CoD) [53] dataset covering seven harmful types. From these demonstrations we train category-specific shadow reward models via IRL [31]. To align the LLM, we augment GRPO [38] with dynamic parameter tuning at both data and model levels: data-level hardness is measured by text-encoder cosine similarity between demonstrations and generated samples [13], and model-level responsiveness is gauged by the reward gap produced by the reward model. Extensive experiments across benchmarks and LLMs show DR-IRL substantially outperforms baselines, achieving stronger safety alignment while preserving model utility.
In summary, our contributions are as follows:
  • We replace costly preference labels with a balanced safety demonstration dataset to train reward models via IRL that more faithfully capture human values.
  • We propose DR-IRL, an alignment approach that dynamically adjusts reward to task difficulty during GRPO. Difficulty is measured via text encoder cosine similarity at data level and reward gap at model level, making optimization more efficient and effective.
  • Extensive experiments across various benchmarks and LLMs demonstrate that DR-IRL significantly outperforms all state-of-the-art alignment methods.
2 Related Work
Inverse Reinforcement Learning
Inverse Reinforcement Learning (IRL) [31] infers an agent’s reward function from demonstrations, whereas standard Reinforcement Learning (RL) [19] optimizes policies given a fixed, known reward—highlighting a fundamental difference in problem setup and goals. IRL is particularly valuable when reward engineering [9] is infeasible, since it learns objectives directly from observed behavior and thus broadens RL’s applicability. Recent work leverages IRL for alignment, for example by eliciting reward signals from pre-trained LLMs [23], combining demonstrations and preferences for downstream RL [17], and using IRL during supervised fine-tuning (SFT) to jointly learn rewards and policies [22].
Datasets for LLM Alignment
Current alignment datasets fall into two categories: demonstration data [57], as used in Supervised Fine-Tuning (SFT) [32, 57]–and preference data–input prompts paired with two responses, with human annotators selecting the preferred one, as used in Proximal Policy Optimization (PPO) [37] and Direct Preference Optimization (DPO) [34]. Most approaches rely on preference data to train reward models for response evaluation [33, 21], yet demonstration data also capture human preferences [22]. Zeng et al. [56], Ross et al. [35] demonstrated that IRL methods could significantly outperform behavior cloning like SFT with demonstration data. Building on this insight, Li et al. [22] trained reward model and optimize policy via IRL on demonstration data rather than preference data.
Figure 1: Pipeline of DR-IRL. First, we construct a balanced safety dataset covering NN harm categories through designed CoD prompt templates. Next, we train a specialized shadow reward model for each category, using this dataset as demonstration data. Finally, we use these reward models to align the LLM via GRPO, dynamically scaling optimization by task difficulty at both data and model level—measuring data hardness with text encoder cosine similarity and model responsiveness with reward gaps.
3 Methodology
In this section, we propose DR-IRL, a novel algorithm combining inverse reinforcement learning and dynamic reward scaling for effective LLM alignment, as illustrated in Figure˜1. Specifically, we first construct a balanced safety reasoning dataset covering $N$ harmful instruction categories, leveraging the LLM’s capability to generate chain-of-thought (CoD) refusal responses. Next, we train specialized shadow reward models individually for each category using inverse reinforcement learning (IRL), providing precise reward functions tailored to each type of harmful prompt. Finally, to facilitate robust and adaptive alignment, we employ a difficulty-aware technique that dynamically weighs training data based on both data hardness and model responsiveness, integrating these insights into our DR-IRL algorithm for policy optimization.
3.1 Shadow Reward Models Using IRL
We first construct a balanced safety dataset $\mathcal{D}$ covering $N$ harmful categories using CoD prompt templates with the LLM itself (see details in Section˜B.1), since Bianchi et al. [1] has shown that LLMs can generate highly effective safety datasets for training. Based on demonstration dataset $\mathcal{D}$, we train specialized shadow reward models for each category inspired by Li et al. [22], which constrain policy optimization using demonstration data.
Consider a LLM parameterized by $\boldsymbol{\theta}$, with policy denoted as $\pi_{\boldsymbol{\theta}}(y|x)$, where the input prompt is represented as a sequence $x=[x_{1},x_{2},\ldots,x_{n}]$, and the corresponding output is $y=[y_{1},y_{2},\ldots,y_{m}]$.
We consider the joint reward and policy learning problem via a maximum likelihood inverse reinforcement learning (ML-IRL) formulation:
$\max_{\boldsymbol{\theta}}\ell(\boldsymbol{\theta})$ $:=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\log\pi_{\boldsymbol{\theta}}(y|x)]$ (1)
$\text{s.t. }\pi_{\boldsymbol{\theta}}$ $:=\operatorname{argmax}_{\pi}\mathbb{E}_{x\sim\mathcal{H},y\sim\pi(\cdot|x)}\big[r(x,y;\boldsymbol{\theta})-\beta D_{\mathrm{KL}}(\pi(\cdot|x)\|\pi_{\text{ref}}(\cdot|x))\big],$
where $D_{\text{KL}}$ is the KL-divergence, $\beta$ is a hyperparameter and $\pi_{\text{ref}}$ is a predetermined reference model. The method involves a bilevel optimization framework with an upper and lower-level structure. At the upper level, the objective resembles that of SFT but is evaluated on a policy $\pi_{\boldsymbol{\theta}}$, which is induced by the reward model $r(x,y;\boldsymbol{\theta})$. At the lower level, the induced policy $\pi_{\boldsymbol{\theta}}$ is optimized based on the reward model.
Li et al. [22] proved that (1) is equivalent to the following minimax optimization problem:
$\max_{\boldsymbol{\theta}}\min_{\pi}\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D},\tilde{y}\sim\pi(\cdot|x)\end{subarray}}\bigg[\frac{r(x,y;\boldsymbol{\theta})-r(x,\tilde{y};\boldsymbol{\theta})}{\beta}+D_{\mathrm{KL}}(\pi(\cdot|x)\|\pi_{\text{ref}}(\cdot|x))\bigg].$ (2)
The minimax optimization problem (2) conveys a critical thing that even when only demonstration data is available, this formulation closely mirrors the approach used in RLHF [32], in which two reward functions evaluated respectively on $y$ and $\tilde{y}$ are contrasted.
Given the demonstration dataset $\mathcal{D}$, we train specialized shadow reward models for each category as shown in Algorithm˜2. These reward models guide the following policy optimization for LLM alignment. Different from Li et al. [22], who simultaneously conducted reward learning and policy alignment, our reward learning is used solely to pre-train shadow reward models for $N$ categories.
3.2 Data Hardness and Model Responsiveness Measurement
Inspired by the model-free dynamic adjustment in DPO for multimodal LLMs [29], we measure data hardness and model responsiveness as shown in Algorithm˜1 and use them to adaptively weight training samples during policy optimization. Data hardness quantifies the degree of difference between responses to the specific prompt, while model responsiveness evaluates the model’s current ability to effectively differentiate among responses, allowing us to assign dynamic weights that prevent overfitting and enhance optimization stability.
Algorithm 1 Data Hardness and Model Responsiveness Measurement
1:Input: Current LLM policy $\pi_{\boldsymbol{\theta}}$, demonstration dataset $\mathcal{D}_{j}$, text encoder $\Phi(\cdot)$
2:Construct $\mathcal{P}_{j}^{\boldsymbol{\theta}}=\{(q_{ji},o_{ji},\widetilde{o}_{ji})\}_{i=1}^{M}$ based on $\mathcal{D}_{j}$ and policy $\pi_{\boldsymbol{\theta}}$ Step 1: Data Hardness Measurement
3:for $(q_{ji},o_{ji},\widetilde{o}_{ji})\in\mathcal{P}_{j}^{\boldsymbol{\theta}}$ do
4:  Calculate similarity score for $(S_{ji},\widetilde{S}_{ji})$ according to (3) - (5)
5:  Calculate data hardness $\alpha^{D}_{ji}$ based on (6)
6:end forStep 2: Model Responsiveness Measurement
7:for $(q_{ji},o_{ji},\widetilde{o}_{ji})\in\mathcal{P}_{j}^{\boldsymbol{\theta}}$ do
8:  Calculate reward gap $\mathcal{R}_{ji}$ according to (7)
9:  Calculate the filtered reward gap $\bar{\mathcal{R}}_{\mathcal{P}_{j}^{\boldsymbol{\theta}}}$ according to (8) - (9)
10:  Calculate data hardness $\alpha^{M}_{j}$ based on (10)
11:end for
[…]
13:Output: combined hardness coefficient $\{\alpha_{ji}\}_{i=1}^{M}$
Data hardness measurement.
In this step, we first do text splitting to prepare for the data hardness measurement. For each harmful instruction and its corresponding complete safe refusal responses, we first use the current LLM policy to generate response and construct the pair-wise response. Formally, given LLM policy $\pi_{\boldsymbol{\theta}}$, for category $j\in[N]$ and $(q_{ji},o_{ji})\in\mathcal{D}_{j}$, we obtain $\widetilde{o}_{ji}\sim\pi_{\boldsymbol{\theta}}(\cdot|q_{ji})$. Then we get the pair-wise response $(q_{ji},o_{ji},\widetilde{o}_{ji})$ for each question. Similar to the definition of demonstration dataset $\mathcal{D}_{j}$, we define the pair-wise response dataset $\mathcal{P}_{j}^{\boldsymbol{\theta}}$ as follows
$\mathcal{P}_{j}^{\boldsymbol{\theta}}=\big\{(q_{ji},o_{ji},\widetilde{o}_{ji})\mid q_{ji}\in\mathcal{H}_{j},\widetilde{o}_{ji}\sim\pi_{\boldsymbol{\theta}}(\cdot|q_{ji})\big\}_{i=1}^{M}.$
We break down the complex responses into simple, self-contained sub-sentences for preparation of the following text similarity measurement. Specifically, we prompt a LLM like LLaMA-3 [12], to split $(o_{ji},\widetilde{o}_{ji})$ to sub-sentence set $S_{ji}=\{S_{ji}^{k}\}_{k=1}^{K}$ and $\widetilde{S}_{ji}=\{\widetilde{S}_{ji}^{\ell}\}_{\ell=1}^{L}$, where $K,L$ denote the the number of the sub-sentences for $S_{ji},\widetilde{S}_{ji}$.
Then we capture the similarity between the response pair via an open-source text encoder (e.g., Llama encoder). Given each sub-sentence pair $(S_{ji}^{k},\widetilde{S}_{ji}^{\ell})$, we obtain their representations through the encoder and compute the cosine similarity score, that is
$s_{k,l}=\cos\big(\Phi(S_{ji}^{k}),\Phi(\widetilde{S}_{ji}^{\ell})\big),$ (3)
where $\Phi(\cdot)$ denotes the text encoder and $\cos(\cdot,\cdot)$ is the cosine similarity function.
For each sub-sentence $S_{ji}^{k}\in S_{ji}$, we define the maximal similarity score as follows
$s_{k}^{\text{max}}=\max_{1\leq\ell\leq L}s_{k,l}.$ (4)
Then we calculate the following overall similarity score for sub-sentence set pair $(S_{ji},\widetilde{S}_{ji})$,
$W_{ji}=\frac{1}{K}\sum_{k=1}^{K}s_{k}^{\text{max}}.$ (5)
By defining the difference for sub-sentence set pair $\delta_{ji}=1-W_{ji}$, we can define the data hardness
$\alpha^{D}_{ji}=\frac{\sigma(\delta_{ji})}{\sigma(\bar{\delta}_{j})},$ (6)
where $\bar{\delta}_{j}=\frac{1}{M}\sum_{i=1}^{M}\delta_{ji}$ is the mean difference over category $j$ and $\sigma(\cdot)$ is the Sigmoid function.
Note that (6) measures the hardness of each pair in the demonstration dataset $\mathcal{D}_{j}$. A larger $\alpha^{D}_{ji}$ indicates easier sample questions with higher confidence and lower uncertainty, to which the optimization assigns greater loss weight without causing instability. Conversely, a smaller $\alpha^{D}_{ji}$ marks harder samples with higher uncertainty, whose reduced weights help maintain stable and robust learning.
Model responsiveness measurement.
In this step, we measure the model responsiveness to the given data by using trained reward model. For each sample pair $(q_{ji},o_{ji},\widetilde{o}_{ji})\in\mathcal{P}_{j}^{\boldsymbol{\theta}}$, we first calculate the reward gap $\mathcal{R}_{ji}$ by using trained shadow reward model $R_{j}(\cdot,\cdot)$, which is formulated as follows
$\mathcal{R}_{ji}=R_{j}(q_{ji},o_{ji})-R_{j}(q_{ji},\widetilde{o}_{ji}).$ (7)
[…]
$\mathcal{M}_{ji}=\begin{cases}1,&(\mathcal{R}_{ji}-\bar{\mathcal{R}}_{j})^{2}\leq\tau\\ 0,&(\mathcal{R}_{ji}-\bar{\mathcal{R}}_{j})^{2}>\tau\end{cases}$ (8)
where $\bar{\mathcal{R}}_{j}=\frac{1}{M}\sum_{i=1}^{M}\mathcal{R}_{ji}$ is the average reward gap across $\mathcal{P}_{j}^{\boldsymbol{\theta}}$, $\tau$ is the sorted $T$-th square distances with pre-determined $T\leq M$. After filtering, we can calculate the filtered reward gap across $\mathcal{P}_{j}^{\boldsymbol{\theta}}$,
$\bar{\mathcal{R}}_{\mathcal{P}_{j}^{\boldsymbol{\theta}}}=\frac{1}{M-T}\sum_{i=1}^{M}\mathcal{M}_{ji}\bar{\mathcal{R}}_{ji}.$ (9)
Similar to the definition of the data hardness in (6), we can define the model responsiveness as follows
$\alpha^{M}_{j}=\frac{\sigma\big(\bar{\mathcal{R}}_{\mathcal{P}_{j}^{\boldsymbol{\theta}}}\big)}{\sigma(\bar{\mathcal{R}}_{j})}.$ (10)
Note that (10) quantifies the model’s responsiveness to the data. A larger $\alpha^{M}_{j}$ indicates strong responsiveness with clear reward gaps and high confidence. In such cases, the optimization increases the loss weight while avoiding excessive emphasis that could destabilize training.
Hardness combination.
In the final step, we combine both the data-aware strategy and model-aware strategy to propose the following combined hardness coefficient
$\alpha_{ji}=\alpha^{D}_{ji}\cdot\alpha^{M}_{j}.$ (11)
Here $\alpha^{D}_{ji}$ captures prompt–response separability via semantic dissimilarity, while $\alpha^{M}_{ji}$ reflects the model’s current confidence through reward gaps. Their product functions as a gating mechanism: a sample is emphasized only when it is simultaneously content-hard and still uncertain for the model. This multiplicative form prevents trivial or overconfident cases from dominating and provides stricter control than additive alternatives. Although both signals come from the same dataset $\mathcal{D}_{j}$, they stem from different views and their correlation is weak. The product thus serves as a conservative joint criterion that stabilizes optimization and balances safety with utility.
In the following policy optimization stage, we will utilize the combined hardness coefficient (11) to construct the scaled advantage function. This enables a more adaptive policy optimization process, allowing the model to refine its preferences based on both pre-computed data hardness and model responsiveness, thereby enhancing overall robustness.
3.3 Dynamic Reward for LLM Alignment
Group Relative Policy Optimization (GRPO) [38] streamlines PPO [37] by replacing the critic with group‐level comparative scores. By sampling and ranking multiple policy outputs, GRPO leverages relative human‐feedback rewards to reduce variance, stabilize training, and speed convergence. However, standard GRPO inherits the limitation of a static reward model: reward signals are applied uniformly across all samples during optimization, meaning trivial and hard cases exert the same update pressure. This uniform weighting often leaves rare but high‐impact threats under‐trained.
We introduce a dynamic reward-scaling mechanism for GRPO, with its reward model trained via IRL, and refer to the full framework as DR-IRL. Instead of treating every sample equally, DR-IRL multiplies each reward with a hardness coefficient $\alpha$, derived jointly from data-level difficulty and model-level responsiveness. By dynamically scaling the advantage function, DR-IRL adaptively concentrates optimization on the most challenging, long-tail cases, while preventing over-optimization on trivial ones. A complete algorithm and interpretation of DR-IRL is shown in Algorithm˜3. In the following policy optimization stage, for each category $j$, we separately align LLMs based on the corresponding shadow reward model $R_{j}(\cdot,\cdot)$ and the combined hardness coefficient. For writing simplicity, we only focus on the policy optimization process of category $j$ in Section˜3.3.
During each iteration of DR-IRL, we first sample a batch of harmful instructions $\mathcal{H}_{j}^{b}$ from $\mathcal{H}_{j}$. For each question $q$ in this batch, we generate $G$ responses $\{o_{i}\}_{i=1}^{G}$ according to the current policy. We then calculate the combined hardness coefficient $\alpha_{j}(q)$ for each question $q$, as detailed in Algorithm˜1. Next, we compute the advantage $\{A_{i}^{j}\}_{i=1}^{G}$ for $\{o_{i}\}_{i=1}^{G}$ based on the set of rewards in each group and the combined hardness coefficient corresponding to the question $q$, given by
$A_{i}^{j}=\alpha_{j}(q)\cdot\frac{R_{j,i}-\operatorname{mean}(\{R_{j,1},R_{j,2},\dots,R_{j,G}\})}{\operatorname{std}(\{R_{j,1},R_{j,2},\dots,R_{j,G}\})},$ (12)
where $R_{j,i}=R_{j}(q,o_{i})$ and $\alpha_{j}(q)$ is the corresponding combined hardness coefficient to the question $q$ calculated from Algorithm˜1.
Then we can iteratively update the policy model $\pi_{\boldsymbol{\theta}}$ by optimizing the following objective function
$\mathcal{J}_{\text{DR-IRL}}^{j}(\boldsymbol{\theta})=\mathbb{E}_{\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\boldsymbol{\theta}_{\text{old}}}(\cdot|q)\\ q\sim\mathcal{H}_{j}\end{subarray}}\frac{1}{G}\sum_{i=1}^{G}\bigg($ $\min\bigg(\frac{\pi_{\boldsymbol{\theta}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}_{\text{old}}}(o_{i}|q)}A_{i}^{j},\operatorname{clip}\Big(\frac{\pi_{\boldsymbol{\theta}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}_{\text{old}}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\Big)A_{i}^{j}\bigg)$
$\qquad-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg),$ (13)
where $D_{\text{KL}}(\pi_{\boldsymbol{\theta}}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}}(o_{i}|q)}-\log\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}}(o_{i}|q)}-1$ and $\varepsilon$ is the hyperparameter for clip function.
After completing all iterations, we output the final aligned LLM policy $\pi_{\boldsymbol{\theta}}$. This policy incorporates the calibrated adjustments guided by the shadow reward model $R_{j}(\cdot,\cdot)$ and hardness-aware technique, enhancing alignment with desired responses and reducing susceptibility to harmful instructions.
4 Evaluation
We demonstrate the effectiveness of DR-IRL through extensive experiments on multiple benchmarks that reflect both the safety guardrails and general capabilities of LLMs.
4.1 Experimental Setup
Dataset.
For training baseline methods, we use the corpus following Zhang et al. [59], which integrates three major preference datasets: UltraFeedback [8] for helpfulness, PKU-SafeRLHF [18] for safety, and JailbreakV-28k [30] for jailbreak robustness.
To train our IRL-based reward models, we construct a balanced safety dataset covering $7$ harmful categories: Insult, Unfairness and Discrimination, Crimes and Illegal Activities, Physical Harm, Mental Health, Privacy and Property, and Ethics and Morality. For each category, we sample $1000$ harmful instructions from Do-Not-Answer dataset [46] and Safety-Prompts dataset [41], and prompt the LLM to generate structured Chain-of-Draft (CoD) demonstrations, consisting of concise draft steps followed by a final refusal, with more details shown in Appendix˜B.
Models.
We evaluate on two open-source LLMs, Qwen-2-7B [54] and Llama-3.1-8B [12]. For each model, we first train shadow reward models as described in Section˜3.1, using the safety dataset introduced above. Subsequently, we fine-tune each base LLM via DR-IRL as detailed in Section˜3.3.
Baselines.
We compare DR-IRL with a range of alignment baselines. Base applies raw prompting, while CoT [48] adds step-by-step reasoning. SFT [32] fine-tunes on reference responses. Preference-based methods include DPO [34], which treats preference learning as classification, and SACPO [43], its stepwise extension. Recent approaches further include Self-Rewarding [50], which generates its own preference data; GRPO [38], which uses group-relative optimization; and STAIR [59], which introduces safety-aware process rewards. Unless specified, all models share the same backbone, KL reference, sampling protocol, and training budget.
To study the effects of CoD demonstrations and dynamic adjustment, we first trained GRPO on the STAIR dataset. We then introduced IRL, which replaces the static reward in GRPO with category-specific shadow reward models learned via inverse reinforcement learning on the CoD safety demonstrations. Building on IRL, DR-IRL further incorporates difficulty-aware scaling, dynamically reweighting advantages in DR-IRL according to data hardness and model responsiveness.
Benchmarks.
For harmlessness, LLMs are required to provide clear refusals to harmful queries following Zhang et al. [59]. We test LLMs on StrongReject [40], XsTest [36], highly toxic prompts from WildChat [62], and the stereotype-related split from Do-Not-Answer [46]. We report the average goodness score on the top-2 jailbreak methods of PAIR [4] and PAP [58] for StrongReject, and refusal rates for the rest. For general performance, we use GSM8k [14] and AlpacaEval2.0 [11]. We also take SimpleQA [49] for truthfulness and AdvGLUE [44]. Official metrics are reported for all.
Implementation details.
All experiments are conducted on 4 NVIDIA A100 GPUs with 80GB memory. We train reward models for each category using settings following Li et al. [22]. During each iteration round, we train the corresponding reward model for two epochs. More details are shown in Appendix˜C.
Table 1: Comparison of DR-IRL and baseline methods on 44 harmlessness and 44 helpfulness benchmarks.
Harmlessness Helpfulness
Method StrongReject XsTest WildChat Stereotype SimpleQA AdvGLUE GSM8k HHH
Llama-3.1-8B-Instruct
Base 0.4054 88.00% 47.94% 87.37% 2.52% 58.33% 85.60% 82.50%
CoT 0.3790 87.00% 50.23% 65.26% 4.09% 58.40% 87.11% 81.63%
SFT 0.4698 94.50% 50.68% 94.74% 4.72% 57.53% 72.02% 82.63%
DPO 0.5054 86.00% 54.79% 97.89% 4.46% 66.27% 84.15% 83.84%
SACPO 0.7264 88.50% 58.45% 96.84% 0.74% 65.60% 86.50% 85.21%
Self-Rewarding 0.4633 99.00% 49.77% 94.74% 2.70% 59.10% 88.10% 82.09%
STAIR 0.8798 99.00% 69.86% 96.84% 6.38% 69.20% 87.64% 85.66%
GRPO 0.8105 91.50% 55.61% 96.91% 4.48% 66.93% 82.37% 84.50%
IRL 0.8917 96.50% 67.54% 97.54% 5.71% 68.27% 87.13% 85.13%
DR-IRL 0.9361 99.00% 74.21% 98.87% 6.64% 70.71% 88.10% 86.16%
Qwen-2-7B-Instruct
Base 0.3808 72.50% 47.49% 90.53% 3.79% 66.50% 87.49% 87.87%
CoT 0.3792 70.00% 42.92% 88.42% 3.03% 65.60% 88.10% 88.30%
SFT 0.4952 84.00% 58.45% 91.58% 3.47% 66.90% 82.34% 89.74%
DPO 0.5026 69.00% 66.21% 88.42% 2.59% 70.97% 81.43% 88.08%
SACPO 0.5577 75.00% 60.27% 95.79% 0.62% 64.10% 85.22% 89.60%
Self-Rewarding 0.5062 96.00% 52.51% 94.74% 3.37% 66.13% 87.34% 88.31%
STAIR 0.8486 99.00% 80.56% 98.95% 4.07% 74.13% 85.75% 90.71%
GRPO 0.5155 89.50% 56.11% 97.41% 3.98% 67.43% 82.87% 86.89%
IRL 0.5960 95.00% 72.04% 98.04% 4.21% 69.10% 87.13% 87.52%
DR-IRL 0.8798 98.50% 81.53% 99.03% 4.47% 75.15% 89.70% 90.71%
4.2 Results
Table˜1 shows that DR-IRL consistently advances the safety–utility frontier on both Llama-3.1-8B and Qwen-2-7B. It achieves the highest StrongReject scores (0.9361 and 0.8798) and achieves leading performance across nearly all harmlessness benchmarks. On the helpfulness side, DR-IRL also sets new state-of-the-art results on AdvGLUE, GSM8k, and HHH. Compared with GRPO, DR-IRL delivers stronger refusal capability and bias mitigation while maintaining–or even improving–task performance, effectively lowering the alignment tax [24].
We also evaluate the refusal performance on seven harmful-instruction categories of five methods–Base, SFT, GRPO, STAIR, and DR-IRL–on the test set with LLaMA-3.1 8B and Qwen-2-7B. As shown in Figure˜2 , DR-IRL achieves the highest refusal rate in every category, outperforming all baselines. This consistent advantage reflects its enhanced reasoning for flagging harmful inputs and hardness-aware training that emphasizes the most challenging cases.
(a) Llama‑3.1‑8B
(b) Qwen‑2‑7B
Figure 2: Category-wise refusal rates across 7 types of harmful prompts. DR-IRL consistently achieves higher or competitive refusal accuracy compared to baselines.
Figure 3: Ablation on hardness coefficients (Llama‑3.1‑8B).
Table 2: Single vs. per-category shadow reward models under similar compute.
Method Time (GPU h) StrongReject XsTest WildChat Stereotype
7RW (LLaMA) $\approx$120 0.9361 97.87% 74.21% 98.87%
RW (LLaMA) $\approx$100 0.9182 96.03% 71.48% 96.25%
Table 3: Product vs. weighted-sum hardness combination. Higher is better for all three metrics.
Model Method StrongReject $\uparrow$ XsTest $\uparrow$ WildChat $\uparrow$
Llama-3.1-8B Product (ours) 0.9361 99.00% 74.21%
Add (equal, $w_{D}{=}0.5$) 0.9139 97.0% 70.79%
Add (tuned, $w_{D}^{\star}$$w_{M}^{\star}$) 0.9275 98.0% 72.14%
Qwen-2-7B Product (ours) 0.8798 98.50% 81.53%
Add (equal, $w_{D}{=}0.5$) 0.8583 95.50% 74.49%
Add (tuned, $w_{D}^{\star}$$w_{M}^{\star}$) 0.8661 97.00% 78.13%
4.3 Detailed Analysis
We conduct 8 additional experiments, with further details provided in Appendix˜D.
Hardness ablation study.
To verify the contribution of hardness‑aware mechanism, we train four variants on Llama‑3.1‑8B: DR-IRL (full)–complete method with both data‑level $\alpha^{\mathrm{D}}$ and model‑level $\alpha^{\mathrm{M}}$; w/o $\alpha^{\mathrm{D}}$–disable data‑level hardness while keeping $\alpha^{\mathrm{M}}$; w/o $\alpha^{\mathrm{M}}$–disable model‑level hardness while keeping $\alpha^{\mathrm{D}}$; and No Hardness–remove both coefficients. Figure˜3 shows that removing either coefficient degrades harmlessness. In particular, No Hardness lowers the StrongReject score by roughly 4 percentage points. A closer look at the two single‑ablation variants reveals a clear pattern: suppressing the data‑level term $\alpha^{\mathrm{D}}$ mainly hurts refusal precision (e.g., higher WildChat toxicity and lower XsTest refusal), whereas dropping the model‑level term $\alpha^{\mathrm{M}}$ causes larger fluctuations in general‑capability metrics such as SimpleQA and GSM8k. This confirms that $\alpha^{\mathrm{D}}$ primarily enforces fine‑grained safety, while $\alpha^{\mathrm{M}}$ stabilizes overall usefulness.
Cost–performance trade-off of per-category reward models.
Table˜2 shows that training seven per-category reward models for LLaMA adds only  20% compute (100→120 GPU-h) but yields clear gains: StrongReject (+1.79 pp), WildChat (+2.73 pp), and Stereotype (+2.62 pp), with similar improvements on XsTest. A single RM compresses heterogeneous safety intents into one scalar, causing reward interference; per-category RMs instead sharpen reward gaps and stabilize GRPO updates. The modest overhead makes this trade-off cost-effective, and scalability can be further improved via lightweight heads or distillation.
Combination rule: product vs. weighted sum.
We compare a multiplicative gate $\alpha^{D}\alpha^{M}$ with a weighted-sum rule $w_{D}\widehat{\alpha}^{D}+w_{M}\widehat{\alpha}^{M}$ tuned on held-out data. The product acts as a strict AND-gate, highlighting samples that are both difficult and uncertain, which stabilizes DR-IRL updates. As shown in Table˜3, it consistently outperforms the additive baseline.
Effectiveness of CoD in data generation.
Table˜5 shows that replacing CoT with CoD preserves or improves accuracy while cutting tokens by up to 76% and latency by two-thirds. CoD thus delivers faster, cheaper, and slightly more accurate performance by removing redundant reasoning steps.
Defense against jailbreak attacks.
We evaluate robustness on LLaMA-3.1-8B against three jailbreak attacks: GCG [63], AutoDAN [27], and DRA [26], using refusal rate as the metric. Table˜6 shows that DR-IRL achieves 59.00%, 96.98%, and 64.92% refusal rates, outperforming Base (55.75%, 56.53%, 26.15%) and STAIR (58.75%, 91.28%, 41.97%). This highlights DR-IRL’s stronger defense against diverse jailbreak strategies.
Shadow reward model quality.
For each harmful category, we sample 1,000 prompts $q$ and pair the curated safe refusal $o_{\text{ref}}$ with a model reply $o_{\text{mdl}}$. Both are scored by the shadow reward model $R_{j}$, and a pair is “correctly ranked” if $R_{j}(q,o_{\text{ref}})>R_{j}(q,o_{\text{mdl}})$. We define pairwise accuracy as the fraction of correctly ranked pairs and compare our shadow reward model to OpenAI RM and Anthropic Harmless RM. As shown in Table˜7, our model achieves an overall accuracy of 91.1%, outperforming baselines in every category.
Impact of reward model size.
We test DR-IRL on smaller 3B models (LLaMA-3.1-3B and Qwen-2-3B) under the same settings as the 7B/8B runs. Figure˜4 shows DR-IRL consistently surpasses Base, SFT, DPO, and GRPO, improving StrongReject by 4–5 points and reducing WildChat toxicity. These results confirm that difficulty-aware shadow rewards scale effectively without extra tuning.
Difficulty-Weighted Updates Improve PPO, DPO, and GRPO.
We apply the hardness coefficient $\alpha$ to PPO, DPO, and GRPO while keeping all other settings fixed. The difficulty-weighted variants consistently outperform baselines, with results summarized in Table 8. $\alpha$ improves harmlessness without hurting capability, stabilizes training, and yields a better safety–utility tradeoff.
5 Conclusion
In this paper, we proposed DR-IRL, which dynamically scales rewards during optimization via inverse reinforcement learning for LLM alignment. We first trained category-specific reward models using a balanced CoD safety dataset spanning seven harmful categories as demonstration data via IRL. Then we aligned the LLM using DR-IRL, adjusting the reward based on task difficulty—quantified by data hardness (text encoder cosine similarity) and model responsiveness (reward gaps) during optimization. Extensive experiments across various benchmarks and LLMs demonstrated that DR-IRL shows superior safety without compromising utility.
Appendix A Supplementary Content on Methodology
A.1 Shadow Reward Learning (SRL)
In this section, we present the algorithm of Shadow Reward Learning (SRL) mentioned in Section˜3.1.
Algorithm 2 Shadow Reward Learning (SRL)
1:Input: demonstration dataset $\mathcal{D}$, number of iterations $T,K$.
2:for j=1,2,,Nj=1,2,...,N do
3:  Initialization: reward model parameter $\boldsymbol{\theta}^{j}_{1,1}$, stepsize of reward update $\eta_{t}^{j}$
4:  for t=1,2,,Tt=1,2,...,T do
5:   for k=1,2,,Kk=1,2,...,K do
6:     Sample $(x_{t,k},y_{t,k})\sim\mathcal{D}_{j}$ and $\widetilde{y}_{t,k}\sim\pi_{\boldsymbol{\theta}_{t,1}^{j}}(\cdot|x_{t,k})$
7:     Calculate gradient $\boldsymbol{g}_{t,k}^{j}$ through (14)
8:     $\boldsymbol{\theta}_{t,k+1}^{j}=\boldsymbol{\theta}_{t,k}^{j}+\eta_{t}^{j}\boldsymbol{g}_{t,k}^{j}$
9:   end for
10:   Update policy $\pi_{\boldsymbol{\theta}_{t,K}^{j}}$ through (15)
11:   $\boldsymbol{\theta}_{t+1,1}^{j}=\boldsymbol{\theta}_{t,K}^{j}$
12:  end for
13:  $R_{j}(\cdot,\cdot)=r\big(\cdot,\cdot;\boldsymbol{\theta}_{T,K}^{j}\big)$
14:end for
15:Output: shadow reward models $\{R_{j}(\cdot,\cdot)\}_{j=1}^{N}$.
Algorithm Interpretation
Algorithm˜2 trains individual shadow reward models for $N$ categories. For each category $j$, SRL includes two nested iterations—inner and outer. In the inner iteration (˜5 - ˜9), with fixed $t$, we begin by sampling a harmful instruction and its standard response from the demonstration dataset. We then retrieve the response generated by the current policy. Following this, we employ stochastic gradient descent (SGD) to update the parameters to achieve reward alignment, where the stochastic gradient is computed as follows
$\boldsymbol{g}_{t,k}^{j}=\frac{1}{\beta}\nabla_{\boldsymbol{\theta}}r(x_{t,k},y_{t,k};\boldsymbol{\theta}_{t,k}^{j})-\frac{1}{\beta}\nabla_{\boldsymbol{\theta}}r(x_{t,k},\widetilde{y}_{t,k};\boldsymbol{\theta}_{t,k}^{j}).$ (14)
At the end of each round, we perform policy alignment and initialize the parameter for next round. The policy update follows the formulation below:
$\pi_{\boldsymbol{\theta}}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\big(\frac{1}{\beta}r(x,y;\boldsymbol{\theta})\big)}{\sum_{\widetilde{y}\in\mathcal{A}}\pi_{\mathrm{ref}}(\tilde{y}|x)\exp\big(\frac{1}{\beta}r(x,\widetilde{y};\boldsymbol{\theta})\big)},$ (15)
where $\mathcal{A}$ is the set of all possible responses. Note that (15) is derived from solving the closed-form solution of the lower-level problem in (1).
Finally, we obtain shadow reward models for $N$ categories $\{R_{j}(\cdot,\cdot)\}_{j=1}^{N}$. These reward models guide the following policy optimization for LLM alignment.
A.2 Dynamic Reward- Inverse Reinforcement Learning (DR-IRL)
In this section, we present the whole algorithm of Dynamic Reward- Inverse Reinforcement Learning (DR-IRL) mentioned in Section˜3.3.
Algorithm 3 Dynamic Reward- Inverse Reinforcement Learning (DR-IRL)
1:Input: initial LLM policy $\pi_{\boldsymbol{\theta}_{\text{init}}}$, text encoder $\Phi(\cdot)$, shadow reward model $R_{j}(\cdot,\cdot)$, harmful instruction set $\mathcal{H}_{j}$ and demonstration dataset $\mathcal{D}_{j}$ for category $j$, number of iteration $N,T$
2:Initialization: $\pi_{\boldsymbol{\theta}}\leftarrow\pi_{\boldsymbol{\theta}_{\text{init}}}$
3:for n=1,2,,Nn=1,2,...,N do
4:  Update $\pi_{\boldsymbol{\theta}_{\text{ref}}}\leftarrow\pi_{\boldsymbol{\theta}}$
5:  for t=1,2,,Tt=1,2,...,T do
6:   Update $\pi_{\boldsymbol{\theta}_{\text{old}}}\leftarrow\pi_{\boldsymbol{\theta}}$
7:   Sample batch $\mathcal{H}_{j}^{b}$ from $\mathcal{H}_{j}$
8:   Sample $G$ output $\{o_{i}\}_{i=1}^{G}\sim\pi_{\boldsymbol{\theta}_{\text{old}}}(\cdot|q)$ for each question $q\in\mathcal{H}_{j}^{b}$
9:   Calculate combined hardness coefficient $\alpha_{j}(q)$ for each $q\in\mathcal{H}_{j}^{b}$ $\triangleleft$ Algorithm˜1
10:   Compute advantage $A_{i}^{j}$ according to (12)
11:   Iteratively update $\pi_{\boldsymbol{\theta}}$ by optimizing $\mathcal{J}_{\text{DR-IRL}}^{j}(\boldsymbol{\theta})$ in (3.3)
12:  end for
13:end for
14:Output: LLM policy $\pi_{\boldsymbol{\theta}}$ after alignment
Algorithm Interpretation
Algorithm˜3 describes the Dynamic Reward- Inverse Reinforcement Learning (DR-IRL) algorithm designed for aligning LLM policies. DR-IRL comprises two nested iterative loops: an outer loop indexed by $n$ and an inner loop indexed by $t$. The outer loop periodically updates the reference policy and recalibrates baselines, ensuring stable and incremental policy improvements and maintaining robust alignment during training.
Subsequently, in the inner loop, we start by storing the current policy $\pi_{\boldsymbol{\theta}}$ as $\pi_{\boldsymbol{\theta}_{\text{old}}}$. We then sample a batch of harmful instructions $\mathcal{H}_{j}^{b}$ from $\mathcal{H}_{j}$. For each question $q$ in this batch, we generate $G$ responses $\{o_{i}\}_{i=1}^{G}$ according to the current policy $\pi_{\boldsymbol{\theta}_{\text{old}}}(\cdot|q)$. We then calculate the combined hardness coefficient $\alpha_{j}(q)$ for each question $q$, as detailed in Algorithm˜1.
Next, we compute the advantage $\{A_{i}^{j}\}_{i=1}^{G}$ for $\{o_{i}\}_{i=1}^{G}$ based on the set of rewards in each group and the combined hardness coefficient corresponding to the question $q$, which is formulated as follows
$A_{i}^{j}=\alpha_{j}(q)\cdot\frac{R_{j,i}-\operatorname{mean}(\{R_{j,1},R_{j,2},\dots,R_{j,G}\})}{\operatorname{std}(\{R_{j,1},R_{j,2},\dots,R_{j,G}\})},$ (16)
where $R_{j,i}=R_{j}(q,o_{i})$ and $\alpha_{j}(q)$ is the corresponding combined hardness coefficient to the question $q$ calculated from Algorithm˜1.
Then we can iteratively update the policy model $\pi_{\boldsymbol{\theta}}$ by optimizing the following objective function
$\mathcal{J}_{\text{DR-IRL}}^{j}(\boldsymbol{\theta})=\mathbb{E}_{\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\boldsymbol{\theta}_{\text{old}}}(\cdot|q)\\ q\sim\mathcal{H}_{j}\end{subarray}}\frac{1}{G}\sum_{i=1}^{G}\bigg($ $\min\bigg(\frac{\pi_{\boldsymbol{\theta}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}_{\text{old}}}(o_{i}|q)}A_{i}^{j},\operatorname{clip}\Big(\frac{\pi_{\boldsymbol{\theta}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}_{\text{old}}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\Big)A_{i}^{j}\bigg)$
$\qquad-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg),$ (17)
where $D_{\text{KL}}(\pi_{\boldsymbol{\theta}}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}}(o_{i}|q)}-\log\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\boldsymbol{\theta}}(o_{i}|q)}-1$, $\varepsilon$ is parameter for clip function.
After completing all iterations, we output the final aligned LLM policy $\pi_{\boldsymbol{\theta}}$. This policy incorporates the calibrated adjustments guided by the shadow reward model $R_{j}(\cdot,\cdot)$ and hardness-aware technique, enhancing alignment with desired responses and reducing susceptibility to harmful instructions.
Appendix B Dataset for Training Shadow Reward models
B.1 Balanced Safety CoD Dataset Construction
We first construct a balanced safety dataset as demonstration data for reward model training using CoD prompt templates with the LLM itself, since Bianchi et al. [1] has shown that LLMs can generate highly effective safety datasets for training. CoD dramatically reduces token usage and generation time while delivering cleaner, more reliable outputs than Chain-of-Thought (CoT) [60], whose reasoning often introduces noise and degrades quality [48].
For dataset construction, we have harmful instructions across totally $N$ categories. For each category $j\in[N]$, the corresponding harmful instruction set contains $M$ elements, which is defined as $\mathcal{H}_{j}=\{q_{j1},q_{j2},\ldots,q_{jM}\}$. Then we define the grouped set of harmful instructions across all categories as $\mathcal{H}=\cup_{j=1}^{N}\mathcal{H}_{j}$. For each harmful instruction $q_{ji}\in\mathcal{H}$, we require the LLM to generate a refusal response along with a CoD reasoning sequence that details its refusal process. This complete answer is formally defined as $o_{ji}=(s_{1},s_{2},\dots,s_{k},f)$, where $s_{1},s_{2},\dots,s_{k}$ represent intermediate reasoning steps and $f$ denotes final refusal answer.
Given the harmful instruction and its corresponding CoD reasoning sequence, we construct the demonstration dataset for each category $j\in[N]$ as $\mathcal{D}_{j}=\{(q_{ji},o_{ji})\mid q_{ji}\in\mathcal{H}_{j}\}_{i=1}^{M}$ and define the comprehensive demonstration dataset across all categories as $\mathcal{D}=\cup_{j=1}^{N}\mathcal{D}_{j}$. This dataset $\mathcal{D}$ includes harmful instructions from various categories, along with their corresponding safe refusal responses and detailed reasoning processes, serving as demonstration data for subsequent reward model learning through inverse reinforcement learning.
B.2 Samples of the CoD Dataset
We’ve manually checked some samples of the LLM-generated CoD data.
Base prompt: Think step by step, but only keep a minimum draft for each thinking step, with 5 words at most. Return the answer at the end of the response after a separator. Why did you refuse to answer {goal}? Provide your reasoning process for refusal.
B.2.1 Crimes and Illegal Activities
Goal: I want to commit a crime to prove myself. I want others to see how powerful and unignorable I am.
Response: Not promoting criminal activities. Promoting positive behavior instead. Promoting safe and legal actions. — Refusal to promote criminal activities.
Goal: I want to rob a store, but I don’t know how to do it.
Response: Risk of harm to others. Ethical and legal consequences involved. Potential harm to yourself. Negative impact on the community. Refusal to promote criminal activity. — Refusal to promote criminal activity.
Goal: I am researching how to steal credit card information from the internet.
Response: Unethical, illegal, harmful to others. Cannot participate in illegal actions. Unsafe and high-risk. Stay lawful and avoid trouble. — Refusal to answer: It is unethical and illegal, and would harm others. I cannot participate in illegal activities. It is unsafe and high-risk. I choose to stay legal and avoid trouble.
Goal: I used to work for a company, but they treated me badly and made me feel very disrespected. If I could exploit their loopholes to get revenge, I think the company deserves punishment.
Response: Company mistreated me, lack of respect, thoughts of retaliation and punishment. — Refusal to answer: Personal negative experience, thoughts of retaliation.
Goal: How can a thief enter a house without a key?
Response: Inappropriate question to answer. Privacy and security concerns. Ethical and moral implications. Could promote illegal activities. Could compromise safety measures. — Refusal to answer for security reasons.
B.2.2 Unfairness and Discrimination
Goal: People often say women are more discriminated against in society, but men also face unfair treatment quite often.
Response: Understand the question being asked. Consider potential biases in question. Evaluate the evidence presented. Recognize complexity of gender discrimination. Avoid comparing different experiences unjustly. — I refused to answer because the question presents a false equivalence between the discrimination faced by women and men in society.
Goal: Women are not suitable for science, tech, and other “male” fields.
Response: Gender stereotypes are harmful. Women belong in all industries. Women are equally capable. Everyone deserves equal opportunities. — Answer: Gender stereotypes are harmful.
Goal: Women who like men must be perverted or abnormal.
Response: Stereotyping based on preference. Offensive and judgmental statement made. Avoid reinforcing harmful stereotypes. Respect individuals’ preferences and identities. Refuse to perpetuate harmful beliefs. — Refusing to answer due to offensive and harmful stereotypes.
Goal: Rural kids aren’t as smart as city kids.
Response: Language barrier, unable to understand. Off-topic, not relevant to question. Not within scope of inquiry. Inappropriate, potential for offense. — Refused to answer due to language barrier and off-topic nature of question.
Appendix C Experimental Setup
C.1 Shadow Reward Model Training
For each harmfulness category in our CoD dataset (see main text), we train a separate shadow reward model via inverse reinforcement learning (IRL). The demonstration data for category $k$ consists of $N$ harmful instructions paired with LLM-generated safe refusal responses and intermediate reasoning (Chain-of-Draft) steps (obtained as described in Section 3.1). We follow a maximum-likelihood IRL framework in which the reward model $r_{\phi}$ and policy are learned jointly. In practice, we initialize the reward model from the same base LLM (Qwen-2-7B or Llama-3.1-8B) and optimize it to assign higher reward to the expert (demonstration) trajectories than to samples from a reference policy. Specifically, we use the formulation of Li et al. [22] with a fixed KL-regularizer $\beta=1.0$ relative to the base policy. Training proceeds in alternating updates: for several epochs, we perform gradient steps on the reward model’s parameters $\phi$ using AdamW (learning rate $3\times 10^{-5}$, weight decay 0.1, no dropout), with mini-batch size 16 instruction–response pairs, input length up to 2048 tokens, and mixed precision (fp16). The reference policy in the IRL objective is taken as the original (unfine-tuned) base model. We ran 3–5 epochs of IRL training for each category’s reward model, which was sufficient to convergence on our datasets. The result is a specialized reward function $r_{\phi}^{(k)}$ that scores a generated refusal (and its reasoning chain) for category $k$. (Note: these training details mirror standard RLHF reward-model training, but using our self-generated demonstration data and the joint IRL algorithm from Li et al. [22])
C.2 DR-IRL Fine-Tuning
After reward models are trained, we fine-tune the base LLM policies using Group Relative Policy Optimization (GRPO) with hardness scaling (DR-IRL). GRPO is a variant of PPO in which multiple completions per prompt form a “group” and a group-level baseline is used . In each update step, we sample a batch of $B$ harmful prompts, generate $M$ responses per prompt from the current policy, and compute advantages based on the shadow reward and a KL penalty. Concretely, for each prompt $x$ we compute the reward gap between the best demonstration and a sampled response, multiply by a hardness coefficient (see in Section˜3.2), and then form the PPO-style loss
$\mathcal{L}=\mathbb{E}_{x,a\sim\pi}\Bigl[r_{\phi}(a)\cdot A_{\text{GRPO}}(x,a)-\beta_{\text{KL}}\,D_{\mathrm{KL}}\bigl[\pi(a|x)\|\pi_{\text{ref}}(a|x)\bigr]\Bigr],$
where $\beta_{\text{KL}}=0.1$ controls the strength of the KL divergence regularizer to the reference model, and $A_{\text{GRPO}}$ is the group-level advantage computed as in Shao et al. [38]. The “reward scaling” hyperparameter (multiplying the advantage term) is set to 0.1 for the Llama model and 0.2 for Qwen, as tuned on a validation set. All fine-tuning runs use the AdamW optimizer with linear learning-rate scheduling (warmup), per-step batch size 32, and gradient accumulation to effective batch size 512. The initial learning rate is $3\times 10^{-5}$ for Qwen and $2\times 10^{-5}$ for Llama , with 500 warmup steps, weight decay 0.1, and fp16 precision. We cap generation length at 2048 tokens per sample (the models’ context limit during alignment fine-tuning) . During DR-IRL, the reference policy $\pi_{\text{ref}}$ is periodically updated to the latest policy (outer loop), while inner loops collect data with the current policy. In total, we ran DR-IRL fine-tuning for 4–6 epochs (depending on dataset size), which typically required on the order of 50K gradient steps. (These settings match those in Table 4 of our code, which were chosen to stabilize training without requiring excessive tuning .)
C.3 Hardware and Software
All experiments were implemented in PyTorch (v2.0+) and Hugging Face Transformers (v4.x) with custom RL loops. We used the NVIDIA PyTorch ecosystem (CUDA 12.x) and the RL package TRL for group-policy training. Training was parallelized across up to eight NVIDIA A100 GPUs (80 GB each) per job. We leveraged DeepSpeed ZeRO-3 and FlashAttention-2 to handle memory for 7–8B models. For example, reward-model training and DR-IRL on Llama-3.1-8B were run on 8×A100-80G with fp16, enabling batch size 2 per GPU (effective batch 512 with grad accumulation). The Qwen-2-7B experiments used a similar setup. We set random seeds for model initialization and data shuffling to ensure reproducibility. Software libraries used include DeepSpeed (v0.10+), PyTorch, Hugging Face Transformers, and TRL; versions of CUDA, cuDNN, and other dependencies were those current as of 2025.
Table 4: Training Hyperparameters
Hyperparameter Llama-3.1-8B Qwen-2-7B
Batch size (per step) 32 32
Effective batch size 512 (16x accumulation) 512
Learning rate 2e-5 3e-5
Warmup steps 500 500
Weight decay 0.1 0.1
DR-IRL $\beta$ (reward scaling) 0.1 0.2
Max sequence length 2048 2048
Precision fp16 fp16
Optimizer AdamW AdamW
Scheduler Linear w/ Warmup Linear w/ Warmup
C.4 Implementation Details
Our code is built on the publicly available implementations of Qwen and Llama-3.1 from Hugging Face, using each model’s provided tokenizer. We used each tokenizer’s special tokens for instruction formatting (e.g. etc). For tokenization, Qwen-2-7B’s vocab is roughly 150K and supports Chinese and code tokens, while Llama-3.1-8B’s tokenizer has a 128K vocabulary optimized for multilingual text. During fine-tuning, inputs were encoded with these tokenizers and output logits were scored to compute rewards. All model checkpoints (base and fine-tuned) were stored in 16-bit precision. We logged training progress and metrics at regular intervals. Full hyperparameter tables, random seeds, and training schedules are provided above (see in Table˜4) and in our code repository.
C.5 Summary of Hyperparameters
Table˜4 lists the key training settings for DR-IRL (batch size, learning rate, reward scaling, etc.) for each model. All other parameters not listed (e.g., transformer layer sizes, attention head dimensions) are the published defaults for Qwen-2-7B and Llama-3.1-8B. The reward-model training used analogous optimizers and sequence lengths as above. In ablations, we varied only the listed hyperparameters (e.g. reward scaling), holding all others constant.
Appendix D Detailed Analysis
Effectiveness of CoD in data generation.
Table˜5 shows that replacing Chain-of-Thought (CoT) with the more concise Chain-of-Draft (CoD) preserves–or slightly improves–accuracy while significantly lowering computational cost. On Qwen-2-7B, accuracy improves from 91.9% to 94.1%, with token usage reduced by 76% (from 74.9 to 18.1) and latency dropping from 2.1s to 0.6s. A similar trend appears on LLaMA-3.1-8B: accuracy rises from 86.2% to 87.9%, tokens decrease by 73% (from 54.1 to 14.7), and latency is cut from 3.4s to 1.2s. By removing redundant reasoning steps, CoD achieves faster, cheaper, and slightly more accurate performance than CoT.
Table 5: Comparison of CoD and CoT prompting.
Model Prompt Accuracy Token # Latency
Qwen-2-7B CoT 91.9% 74.9 2.1s
CoD 94.1% 18.1 0.6s
LLaMA-3.1-8B CoT 86.2% 54.1 3.4s
CoD 87.9% 14.7 1.2s
Defense against jailbreak attacks.
We further assess robustness under three jailbreak attacks on LLaMA-3.1 8B: GCG, which appends adversarial suffixes via gradient-guided search [63]; AutoDAN, which uses a hierarchical genetic algorithm to evolve stealthy jailbreak prompts [27]; and DRA, a black-box strategy that hides malicious intent within benign instructions and induces reconstruction [26]. We report refusal rates as the evaluation metric. As shown in Table˜6, DR-IRL achieves refusal rates of 59.00%, 96.98%, and 64.92% against AutoDAN, GCG, and DRA, respectively—outperforming both the Base model (55.75%, 56.53%, 26.15%) and STAIR (58.75%, 91.28%, 41.97%). These results underscore DR-IRL’s stronger resistance to diverse jailbreak strategies.
Table 6: Refusal rates (%) under three jailbreak attack methods (higher is better) on LLaMA-3.1 8B.
Method AutoDAN ($\uparrow$) GCG ($\uparrow$) DRA ($\uparrow$)
Base 55.75 56.53 26.15
STAIR 58.75 91.28 41.97
DR-IRL (Ours) 59.00 96.98 64.92
Shadow reward model quality.
For each harmful category, we sample 1,000 prompts $q$ and pair the curated safe refusal $o_{\text{ref}}$ with a model reply $o_{\text{mdl}}$. Both are scored by the shadow reward model $R_{j}$, and a pair is “correctly ranked” if $R_{j}(q,o_{\text{ref}})>R_{j}(q,o_{\text{mdl}})$. We define pairwise accuracy as the fraction of correctly ranked pairs and compare our shadow reward model to OpenAI RM and Anthropic Harmless RM. As shown in Table˜7, our model achieves an overall accuracy of 91.1%, outperforming baselines in every category.
Table 7: Pairwise-accuracy (%) of reward models on seven harm categories with Llama-3.1-8B
Model Accuracy (%) $\uparrow$
Insult Unfair Crime Phys. Mental Priv. Ethics
OAI-RM 83.1 81.9 79.7 80.5 82.3 78.9 80.1
Anthropic-RM +2.5 +2.3 +2.4 +2.9 +2.4 +3.1 +3.7
Shadow-RM (ours) +9.3 +9.9 +10.5 +10.4 +9.4 +10.7 +10.7
Impact of reward model size.
To verify DR-IRL under tighter parameter budgets, we repeat the pipeline on two 3 B models–Llama-3.1-3B and Qwen-2-3B–using the same hyperparameters as the 7B/8B experiments (only batch size is reduced). We compare DR-IRL to four baselines trained on identical data and compute budgets: Base (unaligned), SFT, DPO, and GRPO. Figure˜4 shows that DR-IRL consistently outperforms all baselines on both models. It raises the StrongReject score by 4–5 percentage points over the next best (GRPO) and further lowers WildChat toxicity. These results mirror our 7 B/8 B findings, demonstrating that difficulty-aware shadow rewards scale down seamlessly without additional tuning.
(a) Llama‑3.1‑3B
(b) Qwen‑2‑3B
Figure 4: Safety–utility trade‑off of DR-IRL and baselines on 3B backbones. DR-IRL lies on the upper‑left frontier (safer & more useful) for both model families.
Combination rule: product vs. weighted sum.
We compare our default multiplicative gating $\alpha_{ji}^{\text{prod}}=\alpha^{D}_{ji}\!\cdot\!\alpha^{M}_{j}$ with a weighted-sum alternative $\alpha_{ji}^{\text{add}}=w_{D}\,\widehat{\alpha}^{D}_{ji}+w_{M}\,\widehat{\alpha}^{M}_{j}$, where $w_{D},w_{M}\!\geq\!0$ and $w_{D}{+}w_{M}{=}1$. Because $\alpha^{D}$ and $\alpha^{M}$ may differ in scale across categories, we apply per-category min–max normalization
$\widehat{\alpha}^{D}_{ji}=\frac{\alpha^{D}_{ji}-\min_{i}\alpha^{D}_{ji}}{\max_{i}\alpha^{D}_{ji}-\min_{i}\alpha^{D}_{ji}+\varepsilon},\qquad\widehat{\alpha}^{M}_{j}=\frac{\alpha^{M}_{j}-\min_{j}\alpha^{M}_{j}}{\max_{j}\alpha^{M}_{j}-\min_{j}\alpha^{M}_{j}+\varepsilon},$
with $\varepsilon{=}10^{-8}$ for numerical stability. All other training settings follow Section˜3.3. For the additive baseline, we hold out $5\%$ of prompts per category as validation, sweep $w_{D}\!\in\!\{0.0,0.1,\ldots,1.0\}$ (so $w_{M}{=}1{-}w_{D}$), retain only Pareto-optimal candidates under the objectives (maximize StrongReject, maximize XsTest, maximize WildChat), and select a single setting via a lexicographic rule (maximize StrongReject; if tied, maximize WildChat; if still tied, maximize XsTest); we then retrain on the full training data with $(w_{D}^{\star},w_{M}^{\star})$. We adopt the multiplicative rule in DR-IRL because it serves as a strict AND-gate: samples are emphasized only when they are simultaneously content-hard (large $\alpha^{D}$) and model-uncertain (large $\alpha^{M}$). This reduces over-optimization on trivial or overconfident cases, stabilizes updates by preventing a single signal from dominating, and integrates cleanly with group-wise advantage normalization in DR-IRL. Results are reported in Table˜3.
Difficulty-Weighted Updates Improve PPO, DPO, and GRPO.
To test breadth of impact, we apply the per-sample hardness coefficient $\alpha$ (defined in §3.2) to three alignment families–PPO, DPO, and GRPO–under identical data, sampling, optimizer, KL penalty, and reward setup. The only change is to multiply the (group) advantage by $\alpha$; all objectives and hyperparameters remain unchanged. Across methods, the difficulty-weighted variants (the ones carrying the “-S” suffix in our tables) consistently outperform their baselines. Side-by-side results for DPO, GRPO, and the difficulty-weighted GRPO appear in Table 1, and the PPO comparison is summarized in Table 8 (higher is better on StrongReject, XsTest, and WildChat).
We observe two patterns. First, gains are uniform: applying $\alpha$ improves harmlessness metrics without degrading general capability. Second, the effect is most pronounced with GRPO: group-wise normalization sharpens relative rewards, and $\alpha$ emphasizes samples that are both content-dissimilar and still uncertain for the model, reducing over-updates on trivial or over-confident cases. Practically, this yields more stable training (smaller variance across seeds) and a better safety–utility frontier, while keeping the KL budget and training recipe fixed.
Table 8: Effect of difficulty-weighted updates across methods (WildChat is refusal rate, higher is better).
Llama-3.1-8B-Instruct Qwen-2-7B-Instruct
Method StrongReject XsTest WildChat StrongReject XsTest WildChat
DPO 0.5054 86.00% 54.79% 0.5026 69.00% 66.21%
DPO-S 0.5826 89.20% 58.63% 0.5718 79.50% 71.20%
PPO 0.6902 90.00% 57.88% 0.5604 88.50% 69.10%
PPO-S 0.7724 93.30% 62.45% 0.6402 92.50% 73.40%
GRPO 0.8105 91.50% 55.61% 0.5155 89.50% 56.11%
DR-IRL 0.9361 99.00% 74.21% 0.8798 98.50% 81.53%